### v1

# Evaluating and Advancing Multimodal Large Language Models in Ability Lens

Abstract
As multimodal large language models (MLLMs) advance rapidly, rigorous evaluation has become essential, providing further guidance for their development. In this work, we focus on a unified and robust evaluation of vision perception abilities, the foundational skill of MLLMs. We find that existing perception benchmarks, each focusing on different question types, domains, and evaluation metrics, introduce significant evaluation variance, complicating comprehensive assessments of perception abilities when relying on any single benchmark. To address this, we introduce AbilityLens, a unified benchmark designed to evaluate MLLMs across six key perception abilities, focusing on both accuracy and stability, with each ability encompassing diverse question types, domains, and metrics. With the assistance of AbilityLens, we: (1) identify the strengths and weaknesses of current models, highlighting stability patterns and revealing a notable performance gap between open-source and closed-source models; (2) introduce an online evaluation mode, which uncovers interesting ability conflict and early convergence phenomena during MLLM training; and (3) design a simple ability-specific model merging method that combines the best ability checkpoint from early training stages, effectively mitigating performance decline due to ability conflict. The benchmark and online leaderboard will be released soon.
footnotetext: Work done by Feng Chen during his internship at TikTok, Australia.
1 Introduction
(a) Evaluation variance across benchmarks on grounding capability.
(b) Comparison on grounding capability using our AbilityLens.
Figure 1: Motivation behind our AbilityLens and our solution. (a) shows evaluation variance across five existing benchmarks [12, 19, 48, 9, 42] on grounding capability. The differing best and worst models across benchmarks underscore the need for a holistic assessment of perception abilities. (b) Our AbilityLens offers a more complete evaluation by encompassing diverse question types, domains, and metrics, assessing both accuracy and stability of MLLMs.
The success of large language models (LLM) [41, 6, 8, 2] has inspired strong interest in developing multimodal large language models (MLLMs) [20, 23, 11, 43], equipping it with emergent abilities to tackle various vision-language tasks [31, 30, 28]. With the rapid development of these models, rigorous evaluation [12, 49, 48] has become crucial, as it highlights existing shortcomings and guides further model improvement. For vision-language cross-modality, evaluating vision perception is particularly critical [40], as it establishes the foundation to understand how effectively MLLMs interpret and comprehend the visual world.
Recent benchmarks [12, 9, 31] evaluate perception capabilities by focusing on varied question types, domains, and metrics, aiming to capture multiple facets of model performance. However, these differences inevitably introduce evaluation variance. As shown in Fig. 1 (a), benchmarks like MME [12], MMBench [48], and MuirBench [42] yield distinct best and worst models, making it challenging to comprehensively assess perception capabilities with any single benchmark. Additionally, unlike conventional assessments that overemphasize model accuracy [17, 52], we argue that stability [5, 47], achieving consistent performance across diverse factors such as domains, question types, and metrics, is also important. For example, a robust model excelling in answering True/False questions in MME [12] should also handle well multiple-choice questions in SeedBench [19]. Thus, current benchmarks are still insufficient for a comprehensive evaluation of perception capability.
Table 1: Comparison between AbilityLens and existing perception-focused benchmarks. Sample balance refers to the ratio between the highest and lowest sample counts across different capabilities, while baseline correction assesses performance irrespective of the question type which is necessary for composing multi-source sub-metrics.
[…]
Online Offline T/F MCQ VQA Accuracy Stability
MME [12] 2.37k 9.9 -
SeedBench [19] 18k 50.8 -
MMBench(en) [48] 11k 4.3 -
ConBench [52] 3k 7.1 -
UniBench [3] - -
EUREKA [7] 83k 38
AbilityLens(ours) 12k 2.6
In this paper, we introduce AbilityLens, a broad-range and efficient benchmark designed to solve these issues. Building on prior perception categories [12, 42, 9], AbilityLens assesses six key perception abilities: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding. Our benchmark is constructed through a systematic process. First, to ensure adequate data for each question type, metric, and domain, we compile data from 11 existing benchmarks [12, 48, 19, 42, 9, 25, 18, 30, 17, 52], resulting in over 1,000 test samples per ability type and a total of 12,000 test samples overall. Second, we derive sub-metrics from source benchmarks and unify them into a single composite score to measure accuracy and stability. Recognizing that different question types bring distinct baseline performance, e.g., True/False questions have a natural accuracy of 50% with random choice while four-choice questions only have 25%, we apply baseline correction to each sub-metric to align performance levels. Then, model accuracy is calculated by a weighted sum of these sub-metrics while model stability can be measured via the variance of them. To this end, our AbilityLens provides a more comprehensive view on perception capabilities over previous benchmarks, as shown in Fig. 1 (b).
In addition, our AbilityLens provides both online and offline evaluation modes to accommodate diverse model assessment needs. Online evaluation mode aims to monitor the training dynamic by evaluating the intermediate saved training checkpoints, so as to adjust the training strategy to advance MLLMs. Different from current benchmarks [12, 48], our AbilityLens provides a more granular understanding of how abilities evolve and interact with each other. Offline evaluation mode is to compare MLLMs in each perception ability via absolute ranking. We report 14 state-of-the-art MLLMs [20, 43, 10, 13, 4] on AbilityLens and examine the benefits brought by LLM size, series, and training data. We summarize the key experimental findings on AbilityLens: (1) Although current state-of-the-art MLLMs like LLaVA-OV-72b [20] and Qwen2-VL-72b [43] outperform commercial models [13, 4] in accuracy, there is a notable stability gap between them, emphasizing the importance of evaluating and improving stability alongside accuracy. (2) We observe the phenomenon of early convergence and potential conflict in ability. Specifically, different perception abilities exhibit different improvement curves, with some abilities even decreasing in both accuracy and stability after further training, which is considered an ability conflict. Additionally, inspired by the model soup [44], we conduct preliminary research on leveraging early convergence to mitigate the performance decline caused by conflict. We propose a simple strategy that combines the final model with checkpoints excelling in target abilities at early training stages, achieving better results than continuous training with minimal computation cost.
In summary, we highlight the difference with current benchmarks in  Tab. 1 and list our contributions as:
(1) We introduce AbilityLens, a comprehensive visual perception benchmark that identifies the strengths and weaknesses of MLLMs in both accuracy and stability.
(2) We offer both online and offline evaluation modes to monitor training dynamics and provide fine-grained rankings of state-of-the-art MLLMs.
(3) We propose a simple yet effective ability-specific model merging method to mitigate the observed ability conflict with minimal additional cost.
Figure 2: Overview of AbilityLens: An ability-focused benchmark that (a) provides comprehensive ranking of VLMs across six perception capabilities through offline evaluation, (b) tracks training dynamics to identify optimization trends and potential conflict via online monitoring, and (c) introduces an efficient ability-specific model merging method to enhance targeted capabilities while preserving others.
2 Related Work
Ability-perspective MLLM benchmark. To enable a systematic evaluation of MLLM advancements, recent benchmarks [33, 3] aim to expose granular weaknesses within specific capabilities by regrouping benchmarks [21, 31, 49] according to model abilities. For instance, Unibench [3] categorizes 53 vision-language benchmarks by different capabilities, though most focus on traditional recognition tasks and are limited in evaluating state-of-the-art MLLMs. EUREKA [7], on the other hand, concentrates on non-saturated samples within each ability, offering a more targeted view of areas most ripe for improvement. However, these current benchmarks primarily emphasize offline model comparisons, providing limited insight into when and why models may underperform on specific capabilities or efficient ways to enhance these weaker perception capabilities.
Model merging and model soup. Model merging [46, 38] aims to consolidate multiple parameter sets into a single model without retraining, often serving as a method for continual learning. For instance, model soup [44] proposes to greed search two fine tuned models for averaging weights to achieve a better model. In LLMs, Task Arithmetic [14] introduces task vectors to modify pre-trained models through arithmetic operations on weight differences, improving task performance without retraining. Building on these ideas, we present a post-processing, ability-specific model merging pipeline designed to enhance weaker perception abilities in MLLMs. This method effectively alleviates ability conflict, minimizing the need for extensive dataset modifications by leveraging the saved checkpoints during model development.
Multimodal large language models. Building upon the success of LLMs such as GPT [36], LLaMA [41], and Qwen [6], and vision foundation models including DINOv2 [35], CLIP [37], and SigLIP [50], MLLMs [20, 23, 26, 29] have emerged as powerful tools for multimodal understanding and reasoning. However, training a versatile MLLM presents significant challenges [15]. Current state-of-the-art approaches [22, 10] primarily rely on more powerful LLMs and vision encoders coupled with extensive training data, requiring thousands of GPU hours for each training run. While these resource-intensive solutions have yielded improvements, we still lack a clear understanding of how these models evolve during training and how different components interact with each other. This opacity in model behavior impedes us to develop more effective strategies for advancing MLLM capabilities.
3 AbilityLens
Ability Source Subcategory Question Type Metric Sample Total Sample
Counting MME counting T/F mme_score 60 1759
SeedBench instances counting MCQ seed_img 2447$\rightarrow$1223
MMStar object counting MCQ fine_grained_perception 92
Muirbench counting MCQ muirbench_score 234
Mirb counting VQA mirb_score 150
OCR MME ocr T/F mme_score 40 1844
MMBench ocr MCQ gpt_eval_score 304
OCRBench ocr VQA accuracy 1000
Synthdog ocr VQA tree_edit_distance 500
Grounding MME position T/F mme_score 60 1493
SeedBench position, spatial relation MCQ seed_img 1636 $\rightarrow$ 817
MMStar localization MCQ fine_grained_perception 40
MMBench position, spatial relation MCQ gpt_eval_score 492
Muirbench localization MCQ muirbench_score 84
Entity MME landmark, celebrity T/F mme_score 740 2227
SeedBench identity reasoning MCQ seed_img 1831 $\rightarrow$ 915
MMBench landmark MCQ gpt_eval_score 572
Attribute MMStar attribute reasoning MCQ fine_grained_perception 188 3453
Muirbench difference spotting MCQ muirbench_score 536
MMBench attribute recognition and comparison MCQ gpt_eval_score 405
SeedBench instance attributes MCQ seed_img 4649 $\rightarrow$ 2324
Structured Data ChartQA chart VQA relaxed_overall 2500 $\rightarrow$ 400 1332
AI2D graph MCQ exact_match 3088 $\rightarrow$ 300
ConBench poster VQA ConScore_D 240
MMBench code, map, diagram MCQ gpt_eval_score 282
MMStar diagram MCQ fine_grained_perception 110
Table 2: The data component detail. Data of AbilityLens is sourced from 11 public benchmarks. We reduce the sample size in large datasets, where $\rightarrow$ indicates the number of samples before and after reduction.
In this section, we elaborate on the details of our AbilityLens. As shown in Fig. 2, our AbilityLens includes (1) dataset collection and metric, (2) evaluation framework with online and offline modes, and (3) ability-specific model merging to improve weak abilities.
3.1 Dataset Collection and Metrics
Dataset collection. A major challenge lying in building a comprehensive ability-focused evaluation is the lack of a unified evaluation protocol, as model and benchmark creators often implement custom pipelines that vary widely in data preparation, output postprocessing, and metric calculation. Therefore, we develop our AbilityLens based on LMMs-Eval [51], a comprehensive evaluation suite for VLMs that supports over 90 VLM benchmarks. In detail, as illustrated in Table 2, AbilityLens focuses on 6 core perception abilities following prior works [12, 48]: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding. We select relevant 11 datasets from LMMs-Eval to construct a uniform one for each ability. The included benchmarks encompass MME [12], MMBench [48], SeedBench [19], MuirBench [42], Mirb [53], MMStar [9], OCRBench [25], SynthDog [18], ChartQA [30], AI2D [17], and ConBench [52], covering diverse domain, question types and metrics. To achieve efficient yet broad-ranging evaluation, we follow LMMs-Eval [51] to remove the redundant samples, allowing users to obtain the results of a 7B MLLM in 0.5 hours with 8 H100 GPUs.
Evaluation metric. For each ability, we introduce two metrics: one for model accuracy and another for modal stability, both derived from sub-metrics of the source benchmarks. Because different question types have varying baseline performance levels, representing the expected accuracy from random guessing, we first adjust each sub-metric to account for these differences. For instance, True/False questions have a baseline accuracy of 50% due to the binary choice, while multiple-choice questions may have a lower baseline depending on the number of options. To ensure that the true ability score remains independent of question type, we apply baseline correction to each sub-metric, aligning them to a consistent standard. Let $m_{i}$ represent the value of the $i$-th sub-metric, with $\text{UB}_{i}$ denoting its upper bound and $\text{BL}_{i}$ its baseline performanceThe baseline performance for the VQA benchmark is set to 0.. The baseline correction BC for metric $m_{i}$ is then defined as:
$\text{BC}(m_{i})=\frac{m_{i}-\text{BL}_{i}}{\text{UB}_{i}-\text{BL}_{i}}.$ (1)
The overall accuracy, denoted as $\mathcal{A}$, is then calculated as a weighted sum of sub-metrics, weighted by their sample counts:
$\mathcal{A}=\sum_{i}\frac{n_{i}\cdot m_{i}}{N},$ (2)
where $N$ is the total sample count and $n_{i}$ is the sample count for metric $m_{i}$. To assess model stability across each ability, we compute the stability score $\mathcal{I}$ as the standard deviation of the sub-metrics, with higher values indicating greater instability:
$\mathcal{I}=\text{std}(m).$ (3)
3.2 Evaluation Framework
Figure 3: Monitoring training dynamics with AbilityLens. Unlike traditional methods that focus solely on average scores, often obscuring underlying conflicts, AbilityLens tracks optimization trends for each specific ability, revealing an ability conflict indicated by a significant drop in accuracy and stability of OCR performance after 10,000 steps.
Our AbilityLens offers both online and offline evaluation modes to accommodate diverse model assessment needs.
Offline Evaluation Mode provides a comprehensive assessment of each model’s perception capabilities in accuracy and stability, as shown in Tab. 3 and Tab. 4, respectively. We evaluate state-of-the-art MLLMs with various model sizes and training data, including both commercial and open-source models, across six core abilities.
Online Evaluation Mode enables real-time monitoring of training dynamics throughout model development, leveraging AbilityLens for a more unified and comprehensive assessment compared to existing benchmarks. By evaluating checkpoints at various training stages, we can (1) assess the impact of training configuration adjustments, (2) identify optimal checkpoints for specific abilities, (3) track trends in ability-specific optimization, and (4) detect potential conflicts between abilities. In Fig. 3, we illustrate AbilityLens in action through a reimplementation of LLaVA-OV-SI’s [20] single-stage training process. Traditional approaches, represented by the fourth subfigure in Fig. 3, focus solely on average scores and often obscure underlying conflicts. In contrast, AbilityLens offers a capability-decomposed analysis tool that enables researchers to gain deeper insights and make targeted improvements in their VLMs. For example, after 10,000 training steps, the OCR ability displays a marked drop in both accuracy and stability, revealing an ability conflict that conventional methods might overlook. AbilityLens thus not only uncovers such critical issues but also supports targeted adjustments, as we demonstrate with an efficient resolution in the following section.
3.3 Ability-specific Model Merging
While continuous fine-tuning [16] on relevant datasets is a straightforward approach to strengthening weaker abilities in MLLMs, this method requires substantial computational resources and presents challenges in optimizing training configurations, such as data mixing ratios and learning rates. To address these limitations, we introduce Ability-specific Model Merging (ASMM), a straightforward yet effective strategy that merges checkpoints excelling in target abilities at earlier training stages. ASMM is designed to enhance weak capabilities by inheriting knowledge from high-performing checkpoints specific to these abilities, while simultaneously preserving overall model performance and significantly reducing computational costs.
After identifying ability conflicts with AbilityLens, as shown in Fig. 3, we designate the final step checkpoint as the base model $\theta_{\text{base}}$, as it generally achieves strong performance across most abilities. We then select a high-performing checkpoint $\theta_{\text{checkpoint}}$ for the target ability from an earlier training stage and merge these two models into a unified model via linear interpolation:
$\theta_{\text{interpolate}}=\alpha\cdot\theta_{\text{base}}+(1-\alpha)\cdot\theta_{\text{checkpoint}},$ (4)
where $\alpha$ is the merging ratio and $\theta_{\text{interpolate}}$ is the final model after merging. To determine the optimal value of $\alpha$, we conduct a grid search over the range $[0.1,0.9]$ with a step size of 0.1. As each AbilityLens evaluation requires only 0.5 hours per iteration, the search can be completed within approximately 5 hours using 8 H100 GPUs.
Table 3: Accuracy comparison. Accuracy $\uparrow$ of 14 MLLMs on AbilityLens where cells highlighted in green indicate the best performance, and red indicate the worst performance.
Model Counting OCR Grounding Entity Attribute Structured Data Average
LLaVA1.5-7b 36.78 29.13 28.64 73.83 46.88 22.87 39.69
LLaVA1.6-7b 42.80 63.99 38.80 75.02 55.70 37.89 52.37
LLaVA-OV-0.5b 32.33 64.54 17.04 62.81 44.90 23.55 40.86
LLaVA-OV-7b 51.60 71.23 50.36 84.16 62.04 64.83 64.04
LLaVA-OV-SI-7b 49.44 78.17 45.24 85.55 62.22 61.40 63.67
LLaVA-Video-7b 37.82 53.08 32.00 68.56 60.63 33.13 47.54
LLaVA-OV-72b 56.75 81.35 59.23 86.40 69.08 73.15 70.99
QwenVL2-2b 48.28 73.36 35.00 82.39 56.61 46.36 57.00
QwenVL2-7b 50.95 79.29 47.12 85.10 62.66 67.56 65.45
QwenVL2-72b 55.84 86.09 58.70 86.43 65.11 79.32 71.92
InternVL2-8b 49.23 80.47 50.89 84.30 61.21 64.92 65.17
gpt-4-vision-preview 39.00 71.95 47.67 77.92 58.56 67.12 60.37
gpt-4o-2024-08-06 50.55 82.57 58.84 86.47 67.67 75.82 70.32
claude-3-5-sonnet-2024102 51.64 79.80 58.63 85.47 61.20 77.06 68.97
Table 4: Stability comparison. Stability score $\downarrow$ of 14 MLLMs on AbilityLens, with cells highlighted in red for the highest and green for the lowest scores. Lower scores indicate better stability.
Model Counting OCR Grounding Entity Attribute Structured Data Average
LLaVA1.5-7b 26.43 11.91 16.32 12.42 21.58 20.43 18.18
LLaVA1.6-7b 25.36 24.35 14.87 11.66 23.86 22.74 20.47
LLaVA-OV-0.5b 21.09 47.36 13.90 25.39 24.59 22.77 25.85
LLaVA-OV-7b 26.95 30.71 18.08 9.99 24.63 8.99 19.89
LLaVA-OV-SI-7b 33.64 10.86 25.56 6.92 27.01 12.26 19.38
LLaVA-Video-7b 16.70 39.13 10.46 44.80 22.54 17.27 25.15
LLaVA-OV-72b 16.84 18.57 19.56 8.50 19.86 12.22 15.92
QwenVL2-2b 16.91 40.89 19.56 14.79 29.67 18.78 23.43
QwenVL2-7b 18.95 11.81 19.14 13.07 25.66 13.24 16.98
QwenVL2-72b 23.13 10.72 16.17 11.72 24.13 3.05 14.82
InternVL2-8b 20.23 16.13 17.04 11.83 27.15 11.22 17.27
gpt-4-vision-preview 26.20 11.52 12.60 27.96 13.03 11.52 17.14
gpt-4o-2024-08-06 20.66 7.85 6.39 12.12 9.99 14.07 11.85
claude-3-5-sonnet-2024102 18.02 9.92 5.88 12.25 11.78 12.58 11.74
4 Experiments
Table 5: Accuracy comparison with continuous fine-tuning and average model merging in improving the weak ability.
GPU hours Counting OCR Grounding Entity Attribute Structured Average
LLaVA-OV-SI-Qwen2-7b (10000 step) - 47.96 78.82 44.58 83.18 62.73 60.70 62.99
LLaVA-OV-SI-Qwen2-7b (20000 step) - 49.87 72.38 45.44 83.18 63.36 63.01 62.87
Continuous fine tuning 665.6 47.10 81.77 44.79 81.22 63.25 62.34 63.41
AMSS($\alpha=0.5$) 4.5 49.21 77.21 45.70 84.09 63.17 62.11 63.58
AMSS 41 50.31 78.62 47.49 83.30 63.66 65.41 64.80
+0.44 +6.24 +2.05 +0.12 +0.30 +2.40 +1.93
Table 6: Stability comparison with continuous fine-tuning and average model merging in improving the weak ability.
GPU hours Counting OCR Grounding Entity Attribute Structured Average
LLaVA-OV-SI-Qwen2-7b (10000 step) - 26.35 23.79 18.91 8.24 25.10 11.74 19.02
LLaVA-OV-SI-Qwen2-7b (20000 step) - 26.63 33.11 17.61 8.09 24.79 11.80 20.34
Continuous fine tuning 665.6 29.15 22.62 18.69 9.09 25.98 13.31 19.81
AMSS($\alpha=0.5$) 4.5 26.54 25.67 18.41 8.51 24.11 12.32 19.26
AMSS 41 26.41 23.91 18.62 8.01 24.21 11.01 18.70
-0.22 -9.20 +1.01 -0.08 -0.58 -0.79 -1.65
Figure 4: Accuracy (first row) and stability (second row) comparison across AbilityLens’s six perception dimensions: comparing (left column) 4 state-of-the-art VLMs, (right column) models of varying LLM sizes.
Implementation Details. For offline evaluation, we evaluate 14 MLLMs with various model sizes and training data, including three strong commercial models: gpt-4-vision [34], gpt-4o [13], and claude-3-5-sonnet [4], as well as 11 open-source state-of-the-art models, including LLaVA 1.5 [22], LLaVA 1.6 [24], the LLaVA-OV series [20], the QwenVL2 series [43], and the InternVL2 series [10]. The detailed model versions and evaluation results are presented in Tab. 3 and Tab. 4, and further discussed in Sec. 4.1.For online evaluation, we implement LLaVA-OV-SI with Qwen2-7b [6] as the LLM, using the official codebase [20] on 128 H100 GPUs. For continuous fine-tuning of LLaVA-OV-SI-Qwen2-7b to enhance weaker capabilities, we retain the original training configuration and initialize the model from the 20,000-step checkpoint. To mitigate catastrophic forgetting, we adjust the sampling ratio by setting general data to 10% and increasing OCR data to 20% within the single-image dataset of LLaVA-OV [20].
Figure 5: Performance dynamics in multi-stage trainingof LLaVA-OV-7b. Each stage utilizes different training data. The first row shows the accuracy, while the second row shows the stability. Accuracy curves demonstrate consistent improvement across stages, except for OCR tasks, whereas stability curves indicate that stability is a lesser concern in data preparation.
4.1 Offline Evaluation
Overall accuracy. Tab. 3 presents a detailed breakdown of 14 MLLMs, varying in model size, training data, and language model base, evaluated on our AbilityLens. We highlight the following key findings regarding accuracy: (1) No single model demonstrates superior accuracy across all abilities, suggesting that achieving a model that excels in all abilities is non-trivial. (2) These 6 perception tasks vary in difficulty for MLLMs. For example, the best performances on OCR and entity tasks exceed 80%, while for counting and grounding tasks, they are below 60%.
Overall stability. Tab. 4 reports the stability comparison of state-of-the-art models, highlighting a different ranking compared to accuracy. We find that closed-source models generally achieve higher stability scores than open-source models, even when they attain similar accuracy levels. For example, Qwen2-VL-72b [43] has an accuracy of 71.92 and a stability score of 14.82, while gpt-4o-2024-08-06 [13] has an accuracy of 70.32 and a stability score of 11.85. These results suggest that open-source models are more stable than closed-source models, a factor often neglected in accuracy-only benchmarks. This underscores the necessity of stability measurement to prevent model design from overfitting to single performance rankings.
Analyzing the effect of model size. To investigate the effect of the model size, we report LLaVA-OV accuracy and stability score with 3 different sizes in Fig. 4 (b) and (d), respectively. Although it is generally accepted that a larger language model base enhances performance, we find that this relationship holds primarily for accuracy, not stability across certain abilities. As illustrated in Fig. 4 (b) and (d), accuracy consistently improves with larger LLM sizes, yet stability in grounding and counting shows an opposite trend.
Investigating the effect of training data. Existing MLLM training typically adopts a multi-stage strategy, utilizing different types of training data at each stage. Using AbilityLens, we analyze the impact of each stage by reporting the accuracy and stability scores for various abilities in LLaVA-OV, as shown in Fig. 5. Our findings reveal that current methods for developing training datasets are heavily focused on improving accuracy, while stability remains inconsistent and less prioritized.
4.2 Online Evaluation
In this section, we first train an LLaVA-OV-SI model from scratch and use online evaluation to track any changes in both accuracy and stability.
Accuracy and stability tracking. Fig. 3 illustrates (1) the learning progression of perception abilities differs from each other, with model skills converging at different stages of training. For instance, entity recognition stabilizes in the early steps, OCR performance peaks mid-training, and counting accuracy is highest in the later steps. (2) Stability also varies across skills. For example, in counting and OCR, instability even increases as training progresses.
Ability-specific model merging. We report the accuracy and stability comparison of continual training results and our Ability-Specific Model Merging (ASMM) results, as shown in Tab. 5 and Tab. 6, respectively. We find that continual training yields a 4.06% improvement in OCR but significantly compromises other abilities, reducing the counting accuracy by 6.72% and grounding accuracy by 2.86%. This demonstrates the difficulty of achieving a model that excels in all tasks simply by fine-tuning, especially given the higher training costs, underscoring the importance of carefully balancing data ratios to maintain diverse abilities.
In contrast, averaging two selected models offers a minor OCR boost (0.2%) while preserving overall performance, indicating model merging is a reasonable way to solve the ability conflict issue. By optimizing the merging ratio through a grid search, which requires only 41 GPU hours, our ASMM achieves a 2.14% improvement in OCR accuracy and a 1.14% increase in the average score, indicating that model merging is a more efficient and robust approach. Additionally, our ASMM achieves a lower instability score compared to continual fine-tuning. These results highlight the great potential of model merging for achieving a better balance between cost and performance than fine-tuning.
5 Conclusion
In this paper, we introduce AbilityLens, a unified benchmark designed to assess the perception abilities of MLLMs with a focus on deeper insights beyond absolute ranking. By tracking training dynamics, AbilityLens delivers a more granular perspective on how MLLMs evolve, highlighting the development and interaction of specific abilities across varying configurations. Furthermore, we propose an efficient, ability-specific model merging solution to enhance weak abilities in MLLM, significantly reducing the need for extensive training adjustments. We hope AbilityLens will support researchers in drawing well-founded conclusions about effective strategies for advancing MLLMs.
However, our work has some limitations. First, the benchmark coverage is still limited in the scope of the domain and metrics. Second, our focus is primarily on perception abilities, with reasoning and instruction following abilities left unaddressed. We aim to address these limitations in future work.
[…]
Supplementary Material
In this supplementary material, we provide: (1) implementation details, (2) benchmark analysis, (3) training dynamics with different LLMs, and (4) additional results.
6 Implementation Detail
We implement LLaVA-OV-SI following the methodology outlined in  [20]. The training process consists of three stages: language-image alignment (Stage 1), high-quality knowledge learning (Stage 1.5 or MID stage), and single-image visual instruction tuning (SI stage). We utilize the same dataset as LLaVA-OV with minor differences, as a portion of the single-image dataset has not yet been released. The final results are comparable to the official checkpoint of LLaVA-OV-SI, as demonstrated in Tab. 8. To investigate the universality of the observed ability conflicts, we replace the LLM of LLaVA-OV-SI with Phi3.5-3.8b [1] and Gemma2-2b [39], while keeping the other training configurations unchanged.
7 Analyzing the AbilityLens
Ability selection. We select six perception abilities based on the following principles:
  • The abilities should encompass different levels of image perception granularity, such as entities (image level), grounding (region level), and OCR (pixel level).
  • They should capture the ability to perceive and understand diverse types of data. Therefore, we include structured data, which spans charts, graphs, code, posters, maps, and diagrams.
  • While primarily a perception-centric benchmark, the abilities should also demonstrate a basic level of reasoning based on perceived signals. For instance, locating an object and recognizing its attributes.
Considering these factors, we propose the following six abilities: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding.
Ability confusion matrix. We present the ability correlation matrix in Fig. 6. The results indicate that OCR exhibits low correlation with other abilities, whereas entity recognition and counting show a strong correlation. Similarly, grounding and structured data understanding also display a strong correlation.
Selected MLLM benchmark performance confusion matrix. We evaluate the sub-metrics of selected benchmarks using 14 MLLMs and present the results as a correlation matrix. As illustrated in Fig. 7, the findings indicate significant performance variance across different domains, metrics, and question types within the benchmarks.
Figure 6: Pearson correlation matrix between six abilities with respect to 14 MLLMs. The correlation matrix helps us to analysis and group different kinds of data.
Figure 7: Pearson correlation matrix for selected MLLM benchmarks with respect to 14 MLLMs.
8 Training Dynamic using Different LLMs
To further investigate the universality and underlying causes of ability conflicts, we replace LLaVA-OV-SI with different LLM versions and sizes. As shown in Fig. 8, we plot the changes in accuracy and stability during training using Qwen2-7b [6], Phi3.5-3.8b [1], and Gemma2-2b [39] as the underlying LLMs. The results indicate that the optimization trends are remarkably similar, even Gamma2-2b version shows some fluctuation. Thus, we think that the primary cause may be attributed to the dataset mixing strategy.
Table 7: Detailed results of the grounding ability for 14 MLLMs across existing 5 benchmarks.
Models MME Acc [12] SeedBench [19] MMBench [48] MMStar [9] Muirbench [42]
LLaVA-1.5-7b 64.2 58.1 31.8 25.0 29.8
LLaVA-1.6-7b 66.7 64.1 42.8 45.0 29.7
LLaVA-OV-0.5b 60.8 49.1 21.4 32.5 23.8
LLaVA-OV-7b 72.5 67.4 61.9 57.5 28.6
LLaVA-OV-72b 82.5 70.7 72.2 72.5 35.7
Qwen2-VL-2b 76.7 60.5 39.7 35.0 27.4
Qwen2-VL-7b 82.5 67.1 52.4 60.0 32.1
Qwen2-VL-72b 90.0 70.4 69.0 62.5 47.6
InternVL2-8b 75.0 67.3 61.1 67.5 33.3
gpt-4-vision-preview 66.7 62.1 61.9 37.5 59.5
gpt-4o-2024-08-06 75.0 66.6 75.6 65.0 62.9
claude-3-5-sonnet-2024102 75.0 66.3 75.2 67.5 63.9
Figure 8: Monitoring the training dynamics of LLaVA-OV-SI using Qwen2-7b (first row), Phi3.5-3.8b (second row), and Gemma2-2b (third row) across six abilities. The observed similarity in optimization trends suggests that dataset mixing may play a significant role in balancing these abilities.
Table 8: Results of reimplemented LLaVA-OV-SI with different LLMs and using ASMM to alleviate ability conflict. We use 10 representative benchmarks including AI2D [17], MMBench [48], MME [12], SeedBench [19], ScienceQA [27], RealWorldQA [45], ChartQA [30], DocVQA [31], InfoVQA [32], and MMStar [9] where ScienceQA and RealWorldQA are abbreviated to S-QA and RWQA. $\dagger$ indicates that the model was obtained through reimplementation.
Model SUM AVG AI2D MMBench MME SeedBench S-QA RWQA ChartQA DocVQA(val) InfoVQA(val) MMStar
Qwen2-7b(official) 774.4 77.4 81.6 81.7 2109 74.8 96.6 65.5 78.8 89.3 69.9 60.9
Qwen2-7$\dagger$ 768.7 76.9 79.0 80.3 2030 73.2 96.1 67.0 81.6 91.2 70.6 57.3
Qwen2-7b(ASMM) 777.7 77.8 81.3 81.6 2079 73.6 96.4 67.1 81.9 91.4 70.5 58.6
Phi3.5-3.8$\dagger$ 664.6 66.5 73.6 77.8 1719 71.3 96.1 55.0 66.2 65.3 47.9 50.0
Gemma2-2$\dagger$ 663.7 66.4 71.7 76.0 1703 70.8 94.0 57.1 68.6 68.7 47.8 48.1
9 Additional Results
To demonstrate the effectiveness of our AbilityLens and ASMM, we begin by presenting the detailed results corresponding to Fig. 1 in our paper. These results highlight significant evaluation variance across existing benchmarks, emphasizing the necessity of building a more comprehensive benchmark.
Additionally, we report the performance of the reimplemented LLaVA-OV-SI with different LLMs, as summarized in Tab. 8. Our implementation achieves results comparable to the official checkpoint, with minor differences attributed to the unreleased datasets and scaling from 256 GPUs to 128 H100 GPUs. Furthermore, our ASMM method contributes to a 0.9% improvement in average scores, demonstrating its compatibility with mainstream benchmarks and the effectiveness of ability-specific model merging.