### v2

# Evaluating and Advancing Multimodal Large Language Models in Perception Ability Lens

Abstract
As multimodal large language models (MLLMs) advance rapidly, rigorous evaluation has become essential, providing further guidance for their development. In this work, we focus on a unified and robust evaluation of vision perception abilities, the foundational skill of MLLMs. We find that existing perception benchmarks, each focusing on different question types, domains, and evaluation metrics, introduce significant evaluation variance, complicating comprehensive assessments of perception abilities when relying on any single benchmark. To address this, we introduce AbilityLens, a unified benchmark designed to evaluate MLLMs in six key perception abilities (ranging from counting, OCR, to understanding structural data), focusing on both accuracy and stability, with each ability encompassing diverse types of questions, domains, and metrics. With the assistance of AbilityLens, we: (1) identify the strengths and weaknesses of current main-stream MLLMs, highlighting stability patterns and revealing a notable performance gap between state-of-the-art open-source and closed-source models; (2) uncover interesting ability conflict and early convergence phenomena during MLLM training; (3) reveal the primary reason of ability conflict is data mixing ratio and LLM model size; and (4) discuss the effectiveness of some straightforward strategies e.g., fine-tuning and model merging, to solve the ability conflict. The benchmark and online leaderboard is released in https://github.com/Chenfeng1271/AbilityLens.
(a) Evaluation variance across three benchmarks on entity capability.
(b) Comparison on entity capability using our AbilityLens.
Figure 1: Motivation behind our AbilityLens and our solution. (a) shows evaluation variance across three existing benchmarks [12, 18, 40] on entity capability. The differing best and worst models across benchmarks underscore the need for a holistic assessment of perception abilities. (b) Our AbilityLens offers a more complete evaluation by encompassing diverse question types, domains, and metrics, assessing both accuracy and stability of MLLMs. The sizes of closed-source models are artificial estimates.
1 Introduction
The success of large language models (LLM) [36, 5, 8, 1] has inspired strong interest in developing multimodal large language models (MLLMs) [19, 23, 11, 38], equipping it with emergent abilities to tackle various vision-language tasks [29, 28, 26]. The key capability that distinguishes MLLMs from LLMs is their ability to ‘see’, i.e., their perception skills. Thus, evaluating vision perception is particularly critical [35], as it establishes the foundation to understand how effectively MLLMs interpret and comprehend the visual world.
Recent benchmarks [12, 9, 29] evaluate perception capabilities by focusing on varied question types, domains, and metrics, aiming to capture multiple facets of model performance. However, these differences inevitably introduce evaluation variance. As shown in Figure 1 (a), benchmarks like MME [12], MMBench [40], and SeedBench [18] yield distinct best and worst models, making it challenging to comprehensively assess perception capabilities with any single benchmark. Therefore, unlike conventional assessments that overemphasize model accuracy [16, 44], we argue that stability [4, 39], achieving consistent performance across diverse factors such as domains, question types, and metrics, is also important. This metric quantifies the evaluation variance exhibited by models across diverse evaluation scenarios. For example, a robust model excelling in answering True/False questions in MME [12] should also handle well multiple-choice questions in SeedBench [18]. Thus, current benchmarks are still insufficient for a comprehensive evaluation of perception capability.
In this paper, we introduce AbilityLens, a broad-range and efficient benchmark designed to solve these issues. Building on prior perception categories [12, 37, 9], AbilityLens assesses six key perception abilities: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding. Our benchmark is constructed through a systematic process. First, to ensure adequate data for each question type, metric, and domain, we compile data from 11 existing benchmarks [12, 40, 18, 37, 9, 24, 17, 28, 16, 44], resulting in over 1,000 test samples per ability type and a total of 12,000 test samples overall. Second, we derive sub-metrics from source benchmarks and unify them into a single composite score to measure accuracy and stability. Recognizing that different question types bring distinct baseline performance, e.g., True/False questions have a natural accuracy of 50% with random choice while four-choice questions only have 25%, we apply baseline correction to each sub-metric to align performance levels. Then, the model accuracy of each ability is calculated by a weighted sum of these corresponding corrected sub-metrics. Furthermore, we assess the stability of the target model by computing the variance of its z-scores across sub-metrics, which directly reflects its relative performance compared to all candidate models on each sub-metric. To this end, our AbilityLens provides a more comprehensive view on perception capabilities over previous benchmarks, as shown in Figure 1 (b) and Table 1.
Considering our AbilityLens is more general than existing benchmarks, we further use it to monitor the training dynamic by evaluating the intermediate saved training checkpoints, so as to adjust the training strategy to advance MLLMs. Different from current benchmarks [12, 40], our AbilityLens provides a more granular understanding of how abilities evolve and interact with each other. Besides, we also compare mainstream MLLMs in each perception ability via absolute ranking. We report 18 state-of-the-art MLLMs [19, 38, 10, 13, 3] on AbilityLens and examine the benefits brought by LLM size, series, and training data. We summarize the key experimental findings on AbilityLens: (1) Although current state-of-the-art MLLMs like LLaVA-OV-72b [19] and Qwen2VL-72b [38] outperform commercial models [13, 3] in accuracy, there is a notable stability gap between them as shown in Figure 1 (b), emphasizing the importance of evaluating and improving stability alongside accuracy. (2) We observe the phenomenon of early convergence and potential conflict in abilities during model training. Specifically, different perception abilities exhibit different improvement curves, with some abilities meeting performance decline in both accuracy and stability after further training, which is considered an ability conflict. (3) After extensive experiments across various combinations of visual encoders, LLMs, and datasets, we conclude that the data mixing ratio and LLM model size are the primary factors behind the observed ability conflict. Additionally, we conduct a preliminary discussion on leveraging existing fine-tuning and model merging methods to solve the ability conflict.
Table 1: Comparison between AbilityLens and existing perception-focused benchmarks. Sample balance refers to the ratio between the highest and lowest sample counts across different capabilities, while baseline correction assesses performance irrespective of the question type which is necessary for composing multi-source sub-metrics.
[…]
Online Offline T/F MCQ VQA Accuracy Stability
MME [12] 2.37k 9.9 -
SeedBench [18] 18k 50.8 -
MMBench(en) [40] 11k 4.3 -
ConBench [44] 3k 7.1 -
UniBench [2] - -
EUREKA [7] 83k 38
AbilityLens(ours) 12k 2.6
2 Related Work
Perception-based MLLM benchmarks. In parallel with architectural advancements [19, 6], considerable effort has been dedicated to improving benchmarks for MLLMs, which play a crucial role in steering the development of the next generation of models [41]. As a fundamental capability, perception is essential to the performance of MLLMs. To evaluate this skill, benchmarks such as MME [12], MMStar [9], MMBench [40], and SeedBench [18] have been introduced, aiming to capture data from diverse domains and cover a wide range of question types. However, these differences inevitably introduce evaluation variance, motivating the development of a unified dataset that enables a comprehensive assessment of perception abilities.
Ability-perspective MLLM benchmarks. To enable a systematic evaluation of MLLM advancements, recent benchmarks [30, 2] aim to expose granular weaknesses within specific capabilities by regrouping benchmarks [20, 29, 41] according to model abilities. For instance, Unibench [2] categorizes 53 vision-language benchmarks by different capabilities, though most focus on traditional recognition tasks and are limited in evaluating state-of-the-art MLLMs. EUREKA [7], on the other hand, concentrates on non-saturated samples within each ability, offering a more targeted view of areas most ripe for improvement. However, these current benchmarks primarily emphasize offline model comparisons, providing limited insight into when and why models may underperform on specific capabilities or efficient ways to enhance these weaker perception capabilities.
Multimodal large language models. Building upon the success of LLMs such as GPT [33], LLaMA [36], and Qwen [5], and vision foundation models including DINOv2 [32], CLIP [34], and SigLIP [42], MLLMs [19, 23, 25, 27] have emerged as powerful tools for multimodal understanding and reasoning. However, training a versatile MLLM presents significant challenges [15]. Current state-of-the-art approaches [21, 10] primarily rely on more powerful LLMs and vision encoders coupled with extensive training data, requiring thousands of GPU hours for each training run. While these resource-intensive solutions have yielded improvements, we still lack a clear understanding of how these models evolve during training and how different components interact with each other. This opacity in model behavior impedes us to develop more effective strategies for advancing MLLM capabilities.
3 AbilityLens
Ability Source Subcategory Question Type Metric Sample Total Sample
Counting MME counting T/F mme_score 60 1759
SeedBench instances counting MCQ seed_img 2447$\rightarrow$1223
MMStar object counting MCQ fine_grained_perception 92
Muirbench counting MCQ muirbench_score 234
Mirb counting VQA mirb_score 150
OCR MME ocr T/F mme_score 40 1844
MMBench ocr MCQ gpt_eval_score 304
OCRBench ocr VQA accuracy 1000
Synthdog ocr VQA tree_edit_distance 500
Grounding MME position T/F mme_score 60 1493
SeedBench position, spatial relation MCQ seed_img 1636 $\rightarrow$ 817
MMStar localization MCQ fine_grained_perception 40
MMBench position, spatial relation MCQ gpt_eval_score 492
Muirbench localization MCQ muirbench_score 84
Entity MME landmark, celebrity T/F mme_score 740 2227
SeedBench identity reasoning MCQ seed_img 1831 $\rightarrow$ 915
MMBench landmark MCQ gpt_eval_score 572
Attribute MMStar attribute reasoning MCQ fine_grained_perception 188 3453
Muirbench difference spotting MCQ muirbench_score 536
MMBench attribute recognition and comparison MCQ gpt_eval_score 405
SeedBench instance attributes MCQ seed_img 4649 $\rightarrow$ 2324
Structured Data ChartQA chart VQA relaxed_overall 2500 $\rightarrow$ 400 1332
AI2D graph MCQ exact_match 3088 $\rightarrow$ 300
ConBench poster VQA ConScore_D 240
MMBench code, map, diagram MCQ gpt_eval_score 282
MMStar diagram MCQ fine_grained_perception 110
Table 2: The data component detail of AblilityLens. Data of AbilityLens is sourced from 11 public benchmarks. We reduce the sample size in large datasets, where $\rightarrow$ indicates the number of samples before and after reduction. We also provide further decomposition in the supplementary material.
In this section, we elaborate on the details of our AbilityLens. Our AbilityLens includes (1) dataset collection and metrics, and (2) evaluation framework with online and offline modes.
3.1 Dataset Collection and Metrics
Dataset collection. A major challenge lying in building a comprehensive ability-focused evaluation is the lack of a unified evaluation protocol, as model and benchmark creators often implement custom pipelines that vary widely in data preparation, output postprocessing, and metric calculation. Therefore, we develop our AbilityLens based on LMMs-Eval [43], a comprehensive evaluation suite for VLMs that supports over 90 VLM benchmarks. In detail, as illustrated in Table 2, AbilityLens focuses on 6 core perception abilities following prior works [12, 40]: counting, OCR, attribute recognition, entity extraction, grounding, and structural data understanding. We select relevant 11 datasets from LMMs-Eval to construct a uniform one for each ability. The included benchmarks encompass MME [12], MMBench [40], SeedBench [18], MuirBench [37], Mirb [45], MMStar [9], OCRBench [24], SynthDog [17], ChartQA [28], AI2D [16], and ConBench [44], covering diverse domain, question types and metrics. To achieve efficient yet broad-ranging evaluation, we follow LMMs-Eval [43] to remove the redundant samples, allowing users to obtain the results of a 7B MLLM in 0.5 hour with 8 H100 GPUs.
Evaluation metric. For each ability, we define two performance metrics—model accuracy and model stability—each computed from sub-metrics extracted from the source benchmarks. Because different question types have varying baseline performance levels, representing the expected accuracy from random guessing, we first adjust each sub-metric to account for these differences. For instance, True/False questions have a baseline accuracy of 50% due to the binary choice, while multiple-choice questions may have a lower baseline depending on the number of options. To ensure that the true ability score remains independent of question type, we apply baseline correction to each sub-metric, aligning them to a consistent standard. Let $m_{ij}$ represent the result of model $i$ on the $j$-th sub-metric of an ability, with $\text{UB}_{j}$ denoting the upper bound of the metric and $\text{BL}_{j}$ as its baseline performance [^1] The baseline performance for the VQA benchmark is set to 0.. The baseline correction for the result $m_{ij}$ is then defined as:
$\tilde{m}_{ij}=\frac{m_{ij}-\text{BL}_{j}}{\text{UB}_{j}-\text{BL}_{j}}.$ (1)
The accuracy of model $i$ for each ability, denoted as $\mathcal{A}_{i}$, is then calculated as a weighted sum of corresponding corrected sub-metrics $\tilde{m}_{ij}$, weighted by their sample counts:
$\mathcal{A}_{i}=\sum_{j}\frac{n_{j}\cdot\tilde{m}_{ij}}{N},$ (2)
where $N=\sum_{j}{n_{j}}$ is the total sample count and $n_{j}$ is the sample count for the $j$-th metric. To assess the model stability of each ability, we first use z-score $z_{ij}$ to measure the relative performance of model $i$ with other model candidates in each sub-metric:
$z_{ij}=\frac{m_{ij}-\mu_{j}}{\sigma_{j}},$ (3)
where $\mu_{j}$ and $\sigma_{j}$ are the mean and variance of the results of model candidates on this sub-metric, respectively. In this paper, we use 18 MLLMs from different series, sizes, sources, and LLMs as the candidates, capturing a wider spectrum of architectural variations and performance characteristics. Then, we measure the standard deviation of the resulting z-scores as the stability score $\mathcal{I}$ of an ability, with higher values indicating greater instability:
$\mathcal{I}_{i}=\operatorname{std}(\mathbf{z}_{i}).$ (4)
3.2 Evaluation Framework
Figure 2: Monitoring training dynamics with AbilityLens. Unlike traditional methods that focus solely on average scores, often obscuring underlying conflicts, AbilityLens tracks optimization trends for each specific ability, revealing an ability conflict indicated by a significant performance degradation in accuracy and stability of OCR performance after 10,000 steps.
Our AbilityLens offers both online and offline evaluation modes to accommodate diverse model assessment needs.
Offline Evaluation Mode provides a comprehensive assessment of each model’s perception capabilities in accuracy and stability, as shown in Table 3 and 4, respectively. This approach effectively reduces evaluation bias and variance caused by the single benchmark, enabling a more robust and holistic comparison across models. We evaluate state-of-the-art MLLMs with various model sizes and training data, including both commercial and open-source models, across six core abilities.
Online Evaluation Mode enables real-time monitoring of training dynamics throughout model development, leveraging AbilityLens for a more unified and comprehensive assessment compared to existing benchmarks. By evaluating checkpoints at various training stages, we can (1) assess the impact of training configuration adjustments, (2) identify optimal checkpoints for specific abilities, (3) track trends in ability-specific optimization, and (4) detect potential ability conflicts where some abilities meet performance degradation in both accuracy and stability after further training. Addressing these conflicts during training can lead to more balanced and stable multimodal capabilities, ultimately enhancing the overall effectiveness of the models. In Figure 2, we illustrate AbilityLens in action through a reimplementation of LLaVA-OV-SI’s [19] single-stage training process. Traditional approaches, represented by the fourth subfigure in Figure 2, focus solely on average scores and often obscure underlying conflicts. In contrast, AbilityLens offers a capability-decomposed analysis tool that enables researchers to gain deeper insights and make targeted improvements in their VLMs. For example, in the second figure of Figure 2, after 10,000 training steps, the OCR ability displays a marked performance degradation in both accuracy and stability, revealing an ability conflict that conventional methods might overlook. We will discuss using existing fine-tuning and model merging methods to solve this issue in Sec 5.
Table 3: Accuracy comparison. Accuracy $\uparrow$ of 18 MLLMs on AbilityLens where cells highlighted in green indicate the best performance, and red indicate the worst performance.
Model Counting OCR Grounding Entity Attribute Structured Data Average
LLaVA1.5-7b 36.78 29.13 28.64 73.83 46.88 22.87 39.69
LLaVA1.6-7b 42.80 63.99 38.80 75.02 55.70 37.89 52.37
LLaVA-OV-0.5b 32.33 64.54 17.04 62.81 44.90 23.55 40.86
LLaVA-OV-7b 51.60 71.23 50.36 84.16 62.04 64.83 64.04
LLaVA-OV-SI-7b 49.44 78.17 45.24 85.55 62.22 61.40 63.67
LLaVA-Video-7b 37.82 53.08 32.00 68.56 60.63 33.13 47.54
LLaVA-OV-72b 56.75 81.35 59.23 86.40 69.08 73.15 70.99
Qwen2VL-2b 48.28 73.36 35.00 82.39 56.61 46.36 57.00
Qwen2VL-7b 50.95 79.29 47.12 85.10 62.66 67.56 65.45
Qwen2VL-72b 55.84 86.09 58.70 86.43 65.11 79.32 71.92
Qwen2.5VL-3b 50.27 77.49 40.01 82.56 60.70 56.10 61.19
Qwen2.5VL-7b 50.52 80.10 47.71 86.19 62.60 72.44 66.59
Qwen2.5VL-72b 56.53 85.70 59.62 86.82 67.11 79.20 72.49
InternVL2-8b 49.23 80.47 50.89 84.30 61.21 64.92 65.17
InternVL3-8b 51.12 80.02 51.45 84.55 63.46 70.12 66.79
gpt-4-vision-preview 39.00 71.95 47.67 77.92 58.56 67.12 60.37
gpt-4o-2024-08-06 50.55 82.57 58.84 86.47 67.67 75.82 70.32
claude-3-5-sonnet-2024102 51.64 79.80 58.63 85.47 61.20 77.06 68.97
Table 4: Stability comparison. Stability score $\downarrow$ of 18 MLLMs on AbilityLens, with cells highlighted in red for the highest and green for the lowest scores. Lower scores indicate better stability.
Model Counting OCR Grounding Entity Attribute Structured Data Average
LLaVA1.5-7b 72.54 128.40 42.32 110.40 85.30 46.73 80.95
LLaVA1.6-7b 48.00 72.73 31.47 57.29 28.37 43.95 46.96
LLaVA-OV-0.5b 67.07 110.5 60.61 50.00 82.58 73.21 74.00
LLaVA-OV-7b 65.09 62.05 49.28 26.81 72.56 18.62 59.64
LLaVA-OV-SI-7b 115.80 34.17 86.64 42.46 53.20 25.74 88.54
LLaVA-Video-7b 152.20 60.16 78.28 132.30 28.56 79.74 56.43
LLaVA-OV-72b 96.88 57.49 58.85 40.45 42.29 42.63 56.43
Qwen2VL-2b 47.50 89.28 56.53 20.97 64.92 46.80 54.33
Qwen2VL-7b 44.08 45.69 57.59 12.01 46.24 27.55 38.86
Qwen2VL-72b 88.18 24.78 51.52 31.40 49.34 27.82 44.56
Qwen2.5VL-3b 45.16 72.17 60.24 18.88 52.40 47.09 49.32
Qwen2.5VL-7b 45.10 46.70 56.91 14.09 47.28 22.17 38.71
Qwen2.5VL-72b 82.81 29.70 49.62 39.89 43.10 28.33 45.57
InternVL2-8b 44.16 17.34 47.93 9.74 54.60 60.48 38.54
InternVL3-8b 49.12 24.63 44.41 13.82 51.17 62.05 40.87
gpt-4-vision-preview 85.72 77.08 93.53 80.34 52.74 55.78 64.20
gpt-4o-2024-08-06 86.98 24.01 57.06 8.30 12.29 36.48 37.52
claude-3-5-sonnet-2024102 84.57 24.78 59.68 10.85 25.86 27.80 38.92
4 Experiments
Figure 3: Accuracy and stability comparison across AbilityLens’s six perception dimensions, evaluating four state-of-the-art MLLMs alongside models of different LLM scales.
Implementation Details. For offline evaluation, we evaluate 18 MLLMs with various model sizes and training data, including three strong commercial models: gpt-4-vision [31], gpt-4o [13], and claude-3-5-sonnet [3], as well as 15 open-source state-of-the-art models, including LLaVA 1.5 [21], LLaVA 1.6 [22], the LLaVA-OV series [19], the Qwen2VL series [38], Qwen2.5VL series [6] and the InternVL series [10]. The detailed model versions and evaluation results are presented in Table 3 and 4, and further discussed in Sec. 4.1. For online evaluation, we implement LLaVA-OV-SI with Qwen2-7b [5] as the LLM backbone, using the official codebase [19] on 128 H100 GPUs. For continuous fine-tuning of LLaVA-OV-SI-Qwen2-7b to enhance weaker capabilities, we retain the original training configuration and initialize the model from the final 20,000-step checkpoint. To mitigate catastrophic forgetting, we adjust the sampling ratio by setting general data to 10% and increasing OCR data to 20% for the single-image dataset of LLaVA-OV. For using model merging method Task Arithmetic [14] to alleviate ability conflict, we first collect a small calibration set with 3000 OCR samples from the single-image dataset of LLaVA-OV and tune the 1,$0$-step checkpoint on it to find the task arithmetic. Then we add the arithmetic to the final checkpoint.
Figure 4: Performance dynamics in the middle (MID), single-image (SI), and one-vision (OV) stages of LLaVA-OV-7b. Accuracy curves demonstrate consistent improvement across stages, except for OCR tasks, whereas stability curves indicate that stability is less prioritized in data preparation.
4.1 Offline Evaluation
Overall accuracy. Table 3 presents a detailed breakdown of 18 MLLMs, varying in model size, training data, and LLM backbone, evaluated on our AbilityLens. We highlight the following key findings regarding accuracy: (1) No single model demonstrates superior accuracy across all abilities, suggesting that achieving a model that excels in all abilities is non-trivial. (2) These 6 perception tasks vary in difficulty for MLLMs. For example, the best performances on OCR and entity tasks exceed 80%, while for counting and grounding tasks, they are below 60%. It suggests that curriculum learning could be employed to strategically reorder the training sequence, beginning with simpler tasks and progressively advancing to more challenging ones.
Overall stability. Table 4 and 3 (b) report the stability comparison of SOTA models, highlighting a different ranking compared to accuracy. We find that closed-source models generally achieve lower stability scores than open-source models, even attaining similar accuracy levels. For example, Qwen2-VL-72b [38] has an average accuracy score of 71.92 and an average stability score of 44.56, while gpt-4o-2024-08-06 [13] has an accuracy of 70.32 and a stability score of 37.52. These results suggest that closed-source models are more stable than open-source models, a factor often neglected in accuracy-only benchmarks. This underscores the necessity of stability measurement to prevent model design from overfitting single performance rankings.
4.2 Online Evaluation
In this section, we first train an LLaVA-OV-SI model from scratch and use online evaluation to track any changes in both accuracy and stability.
Accuracy and stability tracking. Figure 2 illustrates (1) the learning progression of perception abilities differs from each other, with model skills converging at different stages of training. For instance, entity recognition stabilizes in the early steps, OCR performance peaks mid-training, and counting accuracy is highest in the later steps. (2) Stability also varies across skills. For example, in counting and OCR, instability even increases as training progresses.
5 Discussion and Analysis
Reason behind ability conflict. To investigate the underlying causes of ability conflict, we evaluate various combinations of visual encoders, LLMs, and datasets. Detailed results are provided in Figure 5, 6 below, and Figure 3 of the supplementary material. Notably, the reasons for ability conflict are model size and data mixing, where different visual encoder+LLM combinations exhibit similar performance trends on the same dataset, while variations in the data mixing ratio and different LLM sizes significantly impact performance changes and even alter the timing and number of ability conflict. Specifically, a weaker LLM can result in more ability conflict in training.
Analyzing the effect of model size. To investigate the effect of the model size, we report LLaVA-OV accuracy and stability score with 3 different sizes in Figure 3 (c) and (d), respectively. Although it is generally accepted that a larger language model base enhances performance, we find that this relationship holds primarily for accuracy, not stability across certain abilities where stability in grounding and counting shows an opposite trend.
Figure 5: Monitoring the training dynamics of LLaVA-OV-SI using Qwen2-7b (first row), Phi3.5-3.8b (second row), and Gemma2-2b (third row) across six abilities. We find that model size significantly impacts the occurrence of ability conflict.
Analyzing the effect of model version and visual encoder. We further investigate the influence of visual encoder and LLM in Table 5 and observe an interesting phenomenon: (1) A strong LLM (e.g., Qwen2 vs. Phi3.5) primarily enhances accuracy, but may not be optimal for ensuring the stability of certain abilities. (2) A strong visual encoder (CLIP vs. SigLIP) can boost both accuracy and stability performance.
Table 5: Ablation study on visual encoder and LLM.
Accuracy
Models Counting OCR Grounding Entity Attribute Structured Average
CLIP-L + Qwen2-7B 46.32 69.91 44.29 78.30 60.19 62.22 60.21
Siglip-400M + Qwen2-7B 49.87 72.37 45.44 83.18 63.35 63.01 62.87
Siglip-400M + Phi3.5-3.8B 37.42 56.23 52.97 73.86 53.31 50.47 54.05
Stability
CLIP-L + Qwen2-7B 60.21 65.35 38.80 48.98 47.35 26.70 47.90
Siglip-400M + Qwen2-7B 59.5 60.38 41.20 40.48 43.30 26.35 45.20
Siglip-400M + LPhi3.5-3.8B 86.49 35.04 37.60 36.91 128.45 27.61 58.64
Investigating the effect of training data. Existing MLLM training typically adopts a multi-stage strategy, utilizing different types of training data at each stage. Using AbilityLens, we analyze the impact of each stage by reporting the accuracy and stability scores for various abilities in LLaVA-OV, as shown in Figure 4. Our findings reveal that current methods for developing training datasets are heavily focused on improving accuracy, while stability remains inconsistent and less prioritized.
Simple solutions to alleviate ability conflict. When conflict arises in only one ability during training, approaches such as continuous fine-tuning or model merging can help mitigate the issue. As demonstrated in Tables 6 and 7, Task Arithmetic [14] generally improves both accuracy and stability compared to continuous fine-tuning, which may be sensitive to data ratio variations. However, addressing conflicts across multiple abilities remains an open challenge, as multi-model merging struggles to preserve the optimal performance of each individual model.
Figure 6: Monitoring training dynamics of LLaVA-OV-SI using original data (first row), 50% original data (second row), and original data reducing 50% structured data (third row) across six abilities.
Table 6: Accuracy comparison $\uparrow$ with continuous fine-tuning and model merging methods in improving the weak OCR ability caused by ability conflict. Notably, the final step checkpoint is deemed as the baseline.
GPU hours Counting OCR Grounding Entity Attribute Structured Average
LLaVA-OV-SI-Qwen2-7b (10000 step) - 47.96 78.82 44.58 83.18 62.73 60.70 62.99
LLaVA-OV-SI-Qwen2-7b (final step) - 49.87 72.38 45.44 83.18 63.36 63.01 62.87
Continuous Fine Tuning 665.6 47.10 81.77 44.79 81.22 63.25 62.34 63.41
Task Arithmetic(addition) [14] 13.7 49.60 78.19 45.50 83.46 63.50 63.17 63.90
Table 7: Stability comparison $\downarrow$ with continuous fine-tuning and model merging nethods in improving the weak OCR ability caused by the ability conflict.
GPU hours Counting OCR Grounding Entity Attribute Structured Average
LLaVA-OV-SI-Qwen2-7b (10000 step) - 55.20 36.01 46.50 40.72 38.30 31.99 41.45
LLaVA-OV-SI-Qwen2-7b (final step) - 59.55 60.30 41.22 40.48 43.31 26.35 45.20
Continuous Fine Tuning 665.6 62.19 45.76 43.39 45.60 48.11 24.51 44.93
Task Arithmetic(addition) [14] 13.7 59.19 37.72 46.19 41.96 44.27 26.08 42.57
6 Conclusion
In this paper, we introduce AbilityLens, a unified benchmark designed to assess the perception abilities of MLLMs with a focus on deeper insights beyond absolute ranking. By tracking training dynamics, AbilityLens delivers a more granular perspective on how MLLMs evolve, highlighting the development and interaction of specific abilities across varying configurations. Furthermore, we discuss some simple solutions to enhance weak abilities in MLLM, but they are only applicable in meeting one ability conflict. We hope AbilityLens will support researchers in drawing well-founded conclusions about effective strategies for advancing MLLMs.
However, our work has some limitations. First, the benchmark coverage is still limited in the scope of the domain and metrics. Second, our focus is primarily on perception abilities, with reasoning and instruction following abilities left unaddressed. We aim to address these limitations in future work.
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We claim our MLLM perception benchmark contribution in the abstract.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We claim our limitations in the conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not have theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Our data comes from public dataset and we provide code in the supplementary.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide the data and code in the supplementary and will open-source it to the public.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We elaborate on the dataset detail and model training settings.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We report the performance with trustworthy LLM eval tools.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Our benchmark is efficient with only 4.5 GPU H100 GPU hours.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our benchmark meets NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We claim broader impact in the supplementary.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our benchmark is collected from public datasets, which are free from such cases.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [No]
  59. Justification: Our benchmark is collected from public datasets, which are free from such cases. We will elaborate on it on the Appendix.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: Our benchmark is well documented.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We use LLM for paper writing.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.